a circular coil is receding from a bar magnet wiyh same velocity as that of maget. Will an emf be induced

Grade:9

1 Answers

vikas askiitian expert
13 years ago

case1)

if bar magnet is moving in opposite direction to the motion of coil then relative motion exists ,

relative velocity = v+v=2v

 

case2)

but if both are moving in same direction then there is no relative otion

relative velocity = v-v=0

condition for induced emf is relative motion bw coil and magnet

 so in case 1 emf is induced but in case 2 emf is zero...